A speed skater moving across frictionless ice at 8.2 m/s hits a 5.7 m -wide patch of rough ice. She slows steadily, then continues on at 5.7 m/s
1
Expert's answer
2019-10-04T13:01:31-0400

What is her acceleration on the rough ice?

We use the formula:


"v_f^2-v_i^2=2as"

"5.7^2-8.2^2=2(5.7)a"

"a=-3.0\frac{m}{s^2}"